radical 3- sinx= sinx + radical 12

Answer by lwsshak3(11628)

radical 3- sinx= sinx + radical 12
solve for x
for interval [0,2π]
√3-sinx=sinx+√12
2sinx=√3-√12=√3-√3*√4=√3(1-√4)=√3(1-2)-√3
2sinx=-√3
sinx=-√3/2
x=7π/6,11π/6(in quadrant III and IV where sin<0)